OpenStudy (anonymous):

how do you solve (f(1+4)/f(1+4))-(1/4)(4/f(1))? please explain...

OpenStudy (turingtest):

do we know the function f(x)?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

there was a graph though....

OpenStudy (turingtest):

ok then, without the graph all we can do is say (f(1+4)/f(1+4))-(1/4)(4/f(1))=1-1/f(1) because f(1+4)/f(1+4)=1 and (1/4)(4/f(1)=1/f(1) so 1-1/f(1) is as simplified as we can make this

OpenStudy (anonymous):

ok thanks! and also, how would you do: